OpenStudy (anonymous):

2x-y=-7 4x-y=-4 (–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2)

OpenStudy (anonymous):

(-1.5, 4)

jhonyy9 (jhonyy9):

subtract first from second 4x-y=-4 2x-y=-7 --------- 2x 0 = 3 x=3/2 4(3/2) -y=-4 2*3 -y=-4 6-y=-4 6+4=y y=10 x=3/2 y=10

jhonyy9 (jhonyy9):

second is right (1,5 ,10)

jhonyy9 (jhonyy9):

x=3/2 =1,5